OpenStudy (anonymous):

How to decide whether a taylor polynomial will give you an over or under estimate @ganeshie8

OpenStudy (p0sitr0n):

Error: |E_n(x)| = |f^(n + 1)(z)(x - a)^(n + 1)/(n + 1)!|, where z is in 0 < z < a.

OpenStudy (p0sitr0n):

The remainder is: R_n(x) = f^(n + 1)(z)(x - a)^(n + 1)/(n + 1)!, where z is in 0 < z < a. If R_n(x) > 0, then the polynomial gives an underestimate; if R_n(x) < 0 then the polynomial gives an overestimate.

OpenStudy (p0sitr0n):

What you basically want is to compute the distance between the taylor polynomial and the real function, and then tell whether this distance is positive or negative

OpenStudy (anonymous):

so if you have this

OpenStudy (anonymous):

then you need to evaluate the third degree polynomial at x=.1. which is greater than 0. @P0sitr0n

OpenStudy (p0sitr0n):

compute the taylor polynomial of degree 3. Lets say it gives you r* compute sin(0.1). Lets say it gives you r_. Then compute r*-r_=R if R>0, then its an overestimate, R<0 it is underestimate

OpenStudy (anonymous):

\[P_3(x)=.1-\frac{ .1^3 }{ 3! }+\frac{ .1^5 }{ 5! }=.0998\]

OpenStudy (anonymous):

x=.1, so it would be an overestimate

ganeshie8 (ganeshie8):

looks you have computed 5th degree polynomial

OpenStudy (anonymous):

thats the taylor series for sin(x) third degree

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

@P0sitr0n any guidance

OpenStudy (anonymous):

siths my dood, you are mad smart. u can guide me in the right direction

OpenStudy (anonymous):

if it just asks for \[P_3(.1)\] would you stop when the degree is 3

OpenStudy (anonymous):

so it would be n=1

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

@Zarkon

ganeshie8 (ganeshie8):

you should stop when the exponent is 3 : \[P_3(x)=x-\frac{x^{\color{red}{3}} }{ 3! }\] \[P_3(0.1)=0.1-\frac{0 .1^{\color{red}{3}} }{ 3! }\]

ganeshie8 (ganeshie8):

it is still underestimate http://www.wolframalpha.com/input/?i=%280.1+-+0.1%5E3%2F3%21%29+-+sin%280.1%29

OpenStudy (anonymous):

it is an underestimate because the answer is greater than 0? @ganeshie8

ganeshie8 (ganeshie8):

your estimation of sin(0.1) using 3rd degree taylor polynomial is \[P_3(0.1)=0.1-\frac{0 .1^{\color{red}{3}} }{ 3! } \approx 0.0998333\] actual value of sin(0.1) : \[\sin(0.1)\approx 0.0998334\]

ganeshie8 (ganeshie8):

Clearly your estimation is less than the actual value of sin(0.1) so yours is an underestimate

OpenStudy (anonymous):

while if i were to have this question on a test, and cant use a calculator to see the actual number of sin(.1), how would i reason this?

ganeshie8 (ganeshie8):

i dont remember these sorry @Zarkon @P0sitr0n

OpenStudy (p0sitr0n):

for extremely small values of x, sinx ~ x.

OpenStudy (anonymous):

so the difference between the taylor polynomial of degree n about x=a. and f(a)=R, where if r is negative its an underestimate, positive overestimate